# On the applicability of Kolmogorov’s theory of probability

### v2

@@ 7 Conclusion
It was shown how to construct a mathematically rigorous, non-relativistic quantum theory for an $N$-body system subject to a time-independent scalar potential, which accords with Kolmogorov’s theory of probability. The two main steps were, first, to employ the Born rule for position to construct the probability space (Lem. 1), and, second, to define physically appropriate random variables in relation to the dynamical theory borrowed from QM (Thm. 1).
@@
This approach of constructing a so called hybrid theory can in principle be applied to other quantum-mechanical models, though one does need additional physical arguments to motivate the respective ‘observables’.
@@
With regards to the foundations of mathematical probability theory, it is noteworthy that the equality of the quantum-mechanical and the physically natural ‘Kolmogorovian’ expectation values may be interpreted as evidence for the validity of Kolmogorov’s theory in the quantum domain. [^12] This was already pointed out in [reddigerMadelungPictureFoundation2017]. See also Chap. 2 in [reddigerProbabilisticFoundationNonrelativistic2022]. Thus, Kolmogorov’s axiomatization of probability may indeed be universally applicable. [^13] The physical need for “quantum probability theory” was already put in question by Szabó [szaboCriticalReflectionsQuantum2001], though on the basis of different arguments and a different approach. To my knowledge, the first person to claim the validity of Kolmogorov’s axioms in the quantum domain was Fényes [fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952].
@@
Of course, it is ultimately an empirical question, whether this approach provides a physically viable competitor to QM. In this respect, differences in prediction are to be explored.
Higher order moments of observables other than position generally differ between the two theories. While equality is again obtained for the energy of any stationary state, it may be worthwhile to consider the energy of quantum systems with a simple time dependence.
@@
A physical problem is that testing such systems generally goes beyond the above model, because the radiation, for instance, needs to be included in the dynamics. In turn, the above argument may need to be generalized to systems with spin. Madelung-type reformulations of the Pauli equation are indeed known [takabayasiDescriptionPauliMatter1957, janossyHydrodynamicalModelWave1966], so that a consideration of the corresponding random variables may provide clues to an experiment.
@@
Acknowledgments I would like to thank Bill Poirier, Mike Scherfner, and Max von Renesse for their indirect support of my research efforts. I also thank Jamal Berakdar for helpful comments, Wolfgang Paul for helpful discussion, and Mario Flory for inviting me to a conference, where I held a similarly titled talk. Furthermore, I am indebted to Anhalt UAS for enabling my participation therein.
@@
Appendix: Proof of Thm. 1
1) The assertion follows directly from the invariance of $\operatorname{dom}\hat{H}$ under time evolution (see e.g. Prop. A.1 in [reddigerMathematicalTheoryMadelung2023]).

### v5

@@ 7 Conclusion
It was shown how to construct a mathematically rigorous, non-relativistic quantum theory for an $N$-body system subject to a time-independent scalar potential, which accords with Kolmogorov’s theory of probability. The two main steps were, first, to employ the Born rule for position to construct the probability space (Lem. 1), and, second, to define physically appropriate random variables in relation to the dynamical theory borrowed from quantum mechanics (Thm. 1).
This theory consists of a (time-dependent) probability space, a collection of (time-dependent) random variables, as well as a time evolution law for those two. Mathematical probability theory then dictates how to compute probabilities, probability distributions, expectation values, etc.
In this sense, the reader is provided with the following synopsis of the theory:
  • Time evolution: The Hamiltonian $\hat{H}$ satisfies the assumptions of Thm. 1 and the initial wave function $\Psi_{0}$ is required to be in $\operatorname{dom}\hat{H}$. For any $t\in\mathbb{R}$, $\Psi_{t}$ is then given via Eq. (6.3a).
  • Probability space: At time $t\in\mathbb{R}$ the probability space is $\bigl{(}\mathbb{R}^{3N},\mathcal{B}^{*}(\mathbb{R}^{3N}),\mathbb{P}_{t}\bigr{)}$, as given by Lem. 1. In particular, the probability measure $\mathbb{P}_{t}$ for the positions of the $N$ bodies in $\mathbb{R}^{3N}$ is defined via the probability density $\rho_{t}=\left\lvert\Psi_{t}\right\rvert^{2}$ (cf. Eq. (4.1)).
  • Random variables: The central, physical random variables are the position of the $a$th body $\vec{r}_{a}$, its (current) velocity $\vec{v}_{a}$ as defined in Eq. (6.3b), its stochastic velocity as defined via Eq. (6.4), and the total energy as defined via Eq. (6.3c). Notable random variables, which are derived from the aforementioned ones, are the various angular momenta, the center of mass, and the total momentum (cf. Sec. 5).
The requirement that $\Psi_{0}\in\operatorname{dom}\hat{H}$ constitutes a stronger regularity condition than imposed in the analog quantum-mechanical theory. Furthermore, no physical significance is ascribed to the respective quantum-mechanical operators, which is independent from that of the aforementioned random variables, though they may still be of use in calculating expectation values, etc.
This is a “hybrid theory”, because, philosophically, geometric quantum theory views the time evolution law above as a consequence of the respective Madelung equations. Accordingly and contrary to what was done in this work, the wave functions $\Psi$ are (indirectly) defined in terms of $\rho$ and $\vec{v}$ (cf. [reddigerMathematicalTheoryMadelung2023]).
@@
The presented approach of constructing such a hybrid theory can in principle be applied to other quantum-mechanical models, though one does need additional physical arguments to motivate the random variables of physical interest.
In order to ease further research in this direction, a brief guideline for how to do so is given as follows:
  1. 1)
    Consider the quantum-mechanical analog model in the “position representation”. Define the (time-dependent) probability space via the respective Born rule for position, as done in Sec. 4 and Lem. 1, in particular.
  2. 2)
    For the given Hamiltonian find a (heuristic) Madelung-type reformulation of the respective Schrödinger equation. Geometric quantum theory asserts, that for physically acceptable models such a reformulation always exists (cf. Secs. 3 and 4 in [reddigerMadelungPictureFoundation2017] and Sec. 2.3 in [reddigerProbabilisticFoundationNonrelativistic2022]). This Schrödinger equation will also hold in geometric quantum theory, with the difference that the (initial) wave functions must be elements of the domain of the Hamiltonian (cf. Thm.1.1).
  3. 3)
    As in Sec. 5, employ the analogy to Newtonian (continuum) mechanics to motivate the physically relevant random variables. Those are to replace the respective quantum-mechanical observables. As in Thm. 1, Radon-Nikodým derivatives may be employed to define those random variables in a mathematically rigorous manner (cf. Thm. 1.2) and to relate their expectation values to the quantum-mechanical ones (cf. Thm. 1.3).
@@
With regards to the foundations of mathematical probability theory, it is noteworthy that the equality of the quantum-mechanical and the physically natural ‘Kolmogorovian’ expectation values may be interpreted as evidence for the validity of Kolmogorov’s theory in the quantum domain. This was already shown heuristically in [reddigerMadelungPictureFoundation2017] and Chap. 2 of [reddigerProbabilisticFoundationNonrelativistic2022], yet this work provides a mathematically rigorous argument.
@@
Of course, it is ultimately an empirical question, whether this approach provides a physically viable competitor to quantum mechanics. In this respect, differences in prediction are to be explored.
As illustrated by Ex. 3, the distributions and higher order moments of observables other than position generally differ between the two theories.
@@
A physical problem encountered in Ex. 3 is that testing such systems may go beyond the model employed, because the radiation, for instance, needs to be included in the dynamics. In turn, to make specific experimental predictions, models accounting for spin and time-dependent electromagnetic radiation may need to be considered. Madelung-type reformulations of the Pauli equation are indeed known [takabayasiDescriptionPauliMatter1957, janossyHydrodynamicalModelWave1966], so that a consideration of the corresponding random variables may provide clues to an experiment.
@@
Acknowledgments R. would like to thank Bill Poirier and Mike Scherfner for their indirect support of his research efforts. He also thanks Jamal Berakdar for helpful comments, Wolfgang Paul for helpful discussion on the BCHSH inequality, and Mario Flory for inviting him to a conference, where he held a similarly titled talk. Furthermore, R. is indebted to Anhalt UAS for enabling his participation therein.
@@
Appendix
Proof of Thm. 1
1) The assertion follows directly from the invariance of $\operatorname{dom}\hat{H}$ under time evolution (see for instance Prop. A.1 in [reddigerMathematicalTheoryMadelung2023]).
@@
Proof of Prop. 1
1) For the first statement see Thm. VIII.7 in [reedMethodsModernMathematical1972]. This also implies that Eq. (5.3) holds. The right hand side of Eq. (5.3) is well-defined because of Thm. 1.1.
2) By Thm. 4.8.(i) in [klenkeProbabilityTheoryComprehensive2013], we have $\left\lvert\Psi_{t}\right\rvert^{2}\equiv\left\lvert\Psi_{0}\right\rvert^{2}$ in $L^{1}$. Upon taking the square root, we find that there exists a real-valued function $\alpha\colon t\mapsto\alpha(t)$ such that $\Psi_{t}\equiv e^{-\mathfrak{i}\alpha(t)/\hbar}\Psi_{0}$ in $L^{2}$. Now apply the spectral theorem (Thm. VIII.4 in [reedMethodsModernMathematical1972]) to rewrite $\hat{H}=U^{-1}\hat{\mathcal{E}}U$ for some multiplication operator $\hat{\mathcal{E}}$ on the $L^{2}$-space of some measure space $\left(\Omega,\mathcal{A},\mu\right)$ and for some unitary $U$. From the definition of $\exp(-\mathfrak{i}t\hat{H}/\hbar)$, it follows that
$\left(\exp(-\mathfrak{i}t\hat{\mathcal{E}}/\hbar)-\exp(-\mathfrak{i}\alpha(t)/\hbar)\right)\left(U\Psi_{0}\right)$ (A.1)
vanishes in $L^{2}$ for all $t\in\mathbb{R}$. Now, pick a representative of $U\Psi_{0}$. Then there exists a set of full measure such that for the representative $U\Psi_{0}$ the above expression vanishes pointwise on that set. Since $U\Psi_{0}\neq 0$, there exists an $\omega_{0}\in\Omega$ with $(U\Psi_{0})(\omega_{0})\neq 0$. Writing (^$\bigl{(}\hat{\mathcal{E}}U\Psi_{0}\bigr{)}(\omega_{0})=\mathcal{E}(\omega_{0})\,\left(U\Psi_{0}\right)(\omega_{0})$, it follows that $\mathcal{E}(\omega_{0})\,t\equiv\alpha(t)$. This shows the first assertion. The second one holds by definition of $E$. $\blacksquare$